A navigator on a ship sailing on a course of 330°at 16 knots (nautical miles per hour) observes a lighthouse, due north of the ship. Fifteen minutes later, the lighthouse is due east of the ship. How far is the lighthouse at that instant?
1
Expert's answer
2020-10-05T16:18:54-0400

Solution.

The task is shown in the following picture:



Find the angle α:

"\u03b1=360\u00b0-330\u00b0=30\u00b0."

Find the distance that the ship traveled in 1/4 hour (15 minutes) at a speed 16 nautical miles per hour:

"c=v\cdot t=16\cdot\frac{1}{4}=4" nautical miles.

Find the distance between the ship and the lighthouse:

"b=c\cdot cos(30°)=4\cdot\frac{\sqrt{3}}{2}=2\sqrt{3}\approx3.46" nautical miles.

Answer: the lighthouse is about 3.46 nautical miles from the ship.